Reports until 18:10, Thursday 19 September 2024
Lasers Laser Filter Cavity (GQuEST)
Daniel Grass - posted 18:10, Thursday 19 September 2024 (11872)
Holometer End Cubes designated for LFC Moved to B111B

[Jeff, Ian, Torrey, Alex, Daniel]

We turned off the lasers in B102 for safety purposes (we'd be opening and closing the curtains frequently) and moved both Holometer End Cubes designated for the LFC into B111B onto a 2' tall table.

The next steps are cleaning B111B (with special attention to the mobile clean rooms) and replacing the bottom flange with one with a breadboard pattern.

Images attached to this report